### v1

# From time crystals to time quasicrystals: Exploring novel phases in transverse field Ising chains

Abstract
Time crystals (TCs) and time quasicrystals (TQCs) represent novel phases of matter that arise from the breaking of time translational symmetry in periodically and quasiperiodically driven quantum systems. In this study, we explore the formation of a TQC phase within the disordered quantum Ising chain model under a transverse field (ITF). Notably, TQCs demonstrate robust subharmonic responses at multiple incommensurate frequencies, unlike traditional TCs which respond at a single frequency. Our analysis reveals that the TQC phase exhibits stable magnetization responses even in the presence of interaction perturbations and imperfections in the quasiperiodic driving fields. Employing exact diagonalization techniques, we find that increasing the chain length further stabilizes both TC and TQC phases. These results suggest promising pathways for experimental realization of TQCs in cold atomic systems and quantum simulators, opening avenues for deeper investigation into these intriguing dynamical phenomena.
I Introduction
Time crystals (TCs) represent a fascinating phase of matter that emerges from the spontaneous breaking of time translational symmetry (TTS) in periodically driven quantum systemsr1 ; r2 ; r3 ; r4 ; r5 ; r6 . These systems are typically subjected to external periodic perturbations, allowing them to maintain coherence and exhibit long-lived states characterized by discrete time translational symmetry breaking. In this context, preventing Floquet heating is crucial to avoid a transition to infinite temperature, which would lead to the loss of any temporal orderr7 ; r8 .
In disordered closed quantum systems, many-body localization (MBL) acts as a stabilizing mechanism against Floquet heatingr99 ; r9 . MBL enables the system to remain localized in a specific region of the Hilbert space, thus invalidating the eigenstate thermalization hypothesis (ETH). As a result, the quantum order is preserved over extended time scales, as the system retains its initial state within its eigenstates, shielded from thermalization processesr10 ; r11 ; r12 ; Pr12 ; rr12 ; rp12 .
While periodically driven Floquet systems typically exhibit non-conservation of energy due to continuous energy absorption from external fields, the presence of Floquet MBL alters this dynamic. By breaking the ETH, Floquet MBL prevents the system from reaching thermal equilibrium, leading to the emergence of nonequilibrium phases, including time crystalline phasesr13 ; eqr13 ; r14 ; r15 .
In TC phases, the dynamics of observables in the system are characterized by periodic responses that differ from the driving period. Specifically, for a Hamiltonian that oscillates with a period $T$, the response functions of the system will exhibit oscillations with a period $nT$ (where $n\geq 2$)r8 .
Time quasicrystals (TQCs) represent a novel category of dynamical systems distinguished by their capacity to exhibit subharmonic responses to external driving forces, characterized by multiple incommensurate frequencies. Unlike traditional TCs, which are constrained to a singular frequency response, TQCs arise from quasiperiodic driving of a system where the time-dependent Hamiltonian $H(t)$ incorporates irrationally related frequencies. This leads to quasiperiodic dynamics, resulting in a non-repetitive evolution of the system’s state that densely explores a multi-dimensional toroidal phase space, without revisiting previous configurations. Such quasiperiodicity facilitates intricate dynamical behavior, including complex energy level spacings and dense trajectories in phase spacerr17 ; r31 ; r32 .
In contrast to conventional TCs, TQCs exhibit robust subharmonic responses at a minimum of two incommensurate frequenciesSelf-ordered . Specifically, in the TQC phase, some of the order parameters reveals a stable response characterized by multiple frequencies within the quasiperiodically driven system, a phenomenon referred to as the discrete TQC phase. The dynamics of the TQC phase can be interpreted as trajectories traversing a higher-dimensional space with extended periodicityr34 .
In this study, we explore the spontaneous emergence of TQCs within the framework of the quasiperiodically driven quantum Ising chain in a transverse field (ITF) model. In the context of the ITF model, the TQC phase manifests as a robust response in magnetization that reflects multiple frequencies. This phase is characterized by its remarkable stability and coherence in the presence of interaction perturbations and imperfections in quasiperiodic driving fields.
To date, the TQC phases have been observed in various modelsr17 ; FTQ ; FTS ; rr15 ; r18 . In this work, we examine disordered ITF model, initially analyze it under a periodic driving magnetic field, demonstrating the existence of a TC in this system. We further assess the stability of the TC against perturbations from external interactions and imperfections in rotations, finding it to be robust against such disturbances. In the latter part of our study, we investigate the ITF chain under a sinusoidal driving field, establishing that a disordered ITF chain can also exhibit a TQC phase that remains stable against external interactions and rotational imperfections. Our investigation employs exact diagonalization, and our findings indicate that increasing chain length enhances the stability of both TC and TQC phases. Investigating TQCs paves the way for potential experimental realizations in cold atomic systems, quantum simulators, and other platforms capable of simulating quasiperiodic driving.
This paper is structured as follows: In Section II, we introduce the ITF spin chain, which is periodically driven by an external magnetic field, and analyze its magnetization in various directions, demonstrating the existence of a TC phase. We further investigate the robustness of this TC phase against perturbations in Section III. Section IV provides an overview of the definition and general properties of a TQC phase. In Section V, we incorporate a sinusoidal driving field into the disordered ITF chain, revealing the existence of a TQC phase and exploring its stability under different perturbations. Section VI discusses the emergence of a TQC phase within an ITF chain exhibiting a spin glass phase. Finally, we summarize our findings in Section VII.
II TC in disordered ITF model
Let’s consider a spin-$1/2$ Ising chain subjected to a transverse magnetic field (ITF) and periodically driven by an external magnetic field. The spin-spin interactions in this system are described by the following Hamiltonian:
$H(t)=\begin{cases}\hat{H}_{1}=\sum\limits_{i}\hat{\sigma}_{i}^{x},\quad 0\leq t<t_{1}\\ \hat{H}_{2}=\sum\limits_{i}(-J_{i}\hat{\sigma}_{i}^{z}\hat{\sigma}_{i+1}^{z}+h_{i}\hat{\sigma}_{i}^{x}),\quad t_{1}<t\leq T\end{cases}$ (1)
[…]
The situation is different for finite size chains. The system can tunnel between the two degenerate eigenstates at finite size, and consequently the two lowest lying eigenstates of the system are the following $Z_{2}$ symmetric states: |±$|\pm\rangle=\frac{1}{\sqrt{2}}[|\uparrow\uparrow\dots\uparrow\rangle\pm|\downarrow\downarrow\dots\downarrow\rangle]$. These states are the eigenstates of the Ising parity operator $\hat{P}$ with the eigenvalues ±$\pm 1$, they are long-range correlated with vanishing magnetizationr7 .
At finite temperatures, the Mermin-Wagner theorem asserts that one-dimensional systems with discrete symmetries cannot undergo spontaneous symmetry breaking, as thermal fluctuations in these systems are strong enough to disrupt long-range order, resulting in a finite density of excitationsmermin-wagner ,However, the key point to note is that this theory applies exclusively to equilibrium systems.
In the FM phase, these excitations manifest as domain walls. The energy associated with these domain walls is typically independent of their position along the chain, allowing them to move freely. This mobility of domain walls contributes to the destabilization of any potential long-range order, thus reinforcing the conclusion that spontaneous symmetry breaking cannot be realized in one-dimensional systems at finite temperatures.
The excitation states can be described as superpositions of states featuring finite densities of domain walls distributed across various spatial locations. These delocalized and fluctuating domain walls introduce perturbations that disrupt long-range spin correlations at any finite temperature.
The introduction of disorder, characterized by a non-clean regime with $h_{i}\in[-h,h]$ and $J_{i}\in[J/2,3J/2]$, leads to significant alterations in the system’s physical properties. In the absence of external magnetic fields, the system resides in a FM phase characterized by a non-zero magnetization $m^{z}\neq 0$, as all exchange couplings are positive. Upon the application of a weak magnetic field, specifically when $h\ll J$, domain walls emerge along the chain. For a single domain wall located at position $d$, the excited state of the system is represented as $|d\rangle=|\uparrow\uparrow\dots\uparrow\underline{\downarrow}\uparrow\uparrow\dots\rangle$. In contrast to the clean system, where a thermally activated domain wall can move freely, in the presence of disorder, the domain wall becomes immobilized due to the energy cost associated with its dispersion throughout the chain. As the magnetic field increases, the system can sustain a finite density of domain walls, all of which remain pinned by the disorder. It has been demonstrated that for $h<J$, the system exhibits characteristics of the many-body localized (MBL) phase within the FM phase, which remains stable in the presence of perturbations such as interactionsr23 .
We now examine the potential for spontaneous breaking of time translational symmetry within the phase diagram of the floquet ising Transverse field (FITF) model. The time evolution of the system, as described by equation (1), is governed by the Floquet operator given by:
$\hat{U}_{\rm F}=\exp(-it_{2}\hat{H}_{2})\exp(-it_{1}\hat{H}_{1}).$ (2)
With $t_{1}=\frac{\pi}{2}$, and $t_{2}=1$, the Floquet operator is simplified to:
$\hat{U}_{\rm F}=\exp(-i\hat{H}_{2})\hat{X},~{}~{}~{}\hat{X}=\prod\limits_{i}(-1)^{i}\hat{\sigma}_{i}^{x}.$ (3)
Here, the first term captures the time evolution over the interval $\frac{\pi}{2}<t\leq T$ under the Hamiltonian $\hat{H}_{2}$, while the second term represents a spin rotation about the $x$-axis. Through this formalism, we aim to determine the conditions under which time translational symmetry may undergo spontaneous breaking in the context of the FITF model.
Consider the system initially in a fully polarized state along the $z$ axis, represented as $|\psi_{0}\rangle=\otimes_{i=1}^{L}|\uparrow_{i}\rangle$, where $L$ is the total number of spins. By applying the Floquet time evolution operator to this state, we can derive the system’s state at any time $t$, along with the time dependence of various order parameters. We investigate the time dependence of the order parameters, specifically the magnetizations along the $x$, $y$, and $z$ axes, defined as $m^{x,y,z}=\langle\hat{\sigma}^{x,y,z}\rangle$. In the presence of disorder, it is essential to compute ensemble averages of the magnetizations across various random realizations. By averaging over 150 random realizations, we derive the magnetizations $m^{x,y,z}$ for 160 Floquet cycles. Our findings indicate that when $\bar{J}\gg\bar{h}$, the magnetization $m^{z}$ exhibits oscillatory behavior around zero with a period of $2T$ (refer to Fig. 1-(up)). The periodic oscillations of $m^{z}$, with a period that is double that of the Hamiltonian, suggest a breakdown of time translational symmetry. This phenomenon is indicative of time crystal (TC), provided that the observed oscillations remain stable in the thermodynamic limit ($L\rightarrow\infty$) and are resilient against perturbations such as interactions.
Figure 1: Time evolution of the magnetizations along $x,y$, and $z$ directions for a chain of 12 sites. For $J=5.5$ and $h=0.3$ (up), the magnetization along the $z$ axis exhibits periodic oscillations with a period of $2T$, while the $x$ and $y$ components remain zero. The inset plot illustrates the fast Fourier transform (FFT) of $m^{z}(t)$, highlighting a pronounced peak at $\omega=\pi$, which signifies the periodic nature of the magnetization. Conversely, for $J=0.3$ and $h=4.5$ (down), the magnetization oscillations are characterized by a lack of periodicity and stability. The peak value in the FFT is multiplied by 0.01, and this adjustment is applied to all other plots as well.
We also derive the time dependence of all components of the magnetization in the PM phase, where $\bar{J}<\bar{h}$. Our findings, presented in Fig. 1-(down), indicate that while the magnetizations are nonzero, their oscillations exhibit a lack of periodicity and stability.
III Stability of the TC phase
In this section, we investigate the stability of the TC in the presence of various external perturbations. We will demonstrate that the introduction of diverse perturbations, including interactions and imperfect rotations (deviations from perfect rotation), does not compromise the time crystalline order.
III.1 Stability of the TC phase under interactions
In this part, we examine the stability of the TC phase under the influence of exchange interactions along the $x$-axis. In this context, the second term of the Hamiltonian in Eq. (1) is modified as follows:
$\hat{H}_{2}=\sum\limits_{i}(-J_{i}\hat{\sigma}_{i}^{z}\hat{\sigma}_{i+1}^{z}+h_{i}\hat{\sigma}_{i}^{x})+\lambda\sum_{i}\hat{\sigma}_{i}^{x}\hat{\sigma}_{i+1}^{x},$ (4)
where $\lambda$ denotes the strength of the interaction between nearest neighboring spins located at sites $i$ and $i+1$.
To elucidate the origin of the interaction represented by the $\lambda$-term, we can map the spin Hamiltonian (4) onto a fermionic Hamiltonian. By applying the transformation $\sigma_{i}^{x}\leftrightarrow\sigma_{i}^{z}$ and utilizing the Jordan-Wigner (JW) spin-fermion transformationsMB18 , we observe that the $\lambda$-term transforms into $\lambda c_{i}^{\dagger}c_{i}c_{i+1}^{\dagger}c_{i+1}$, where $c_{i}$ ($c_{i}^{\dagger}$) are the annihilation (creation) operators for fermions at site $i$. This expression clearly indicates the interaction between two fermions located at sites $i$ and $i+1$.
[…]
In Fig. 2, we present the time evolution of the magnetization $m^{z}$ in the ferromagnetic many-body localized (FM-MBL) phase for varying interaction strengths. The observed time dependence of $m^{z}$ reveals significant oscillatory behavior, characterized by a period of $2T$, superimposed on an oscillatory-decaying envelope that is dependent on the interaction strength $\lambda$. Specifically, for a finite-size system, the envelope exhibits a prolonged time period at low interaction strengths, and diminishes as $\lambda$ increases. Furthermore, we have computed the Fourier transform of $m^{z}$ as a function of frequency for various $\lambda$ values. For small values of $\lambda$, during the observed time window ($t/T\in[0,160]$), the magnetization envelope curve remains smooth without any knots, and the FFT reveals a distinct subharmonic peak at $\omega=\pi$. As $\lambda$ increases, this peak begins to split due to the interference of magnetization oscillations and eventually disappears, indicating the complete destruction of the TC phase. Extending the time window for analyzing the system’s magnetization increases the probability of knot formation, leading to a double peak in the FFT even for small interaction values.
The observed oscillatory-decaying envelope in the magnetization can be attributed to finite-size effects. To substantiate this assertion, we present in Fig. 3 the time evolution of the magnetization $m^{z}$ at a coupling strength $\lambda<J$ for chains of varying lengths. Notably, as the chain length increases, the period of the envelope oscillations also increases, suggesting that in the limit of infinite chain length ($L\rightarrow\infty$), the envelope disappears, leaving only stable oscillations with a period of $2T$. This behavior indicates the existence of a robust TC for interaction strengths below a critical value $\lambda_{c}$, with $\lambda_{c}$ being dependent on the system size. Furthermore, in the inset of Fig. 3, we present the FFT of $m^{z}$ for different system lengths. For smaller systems, the FFT reveals a double peak structure, reflecting the beating of magnetization oscillations. As the system size increases, these double peaks converge into a single peak at frequency $\omega=\pi$. Continued increases in system size result in a sharper peak that becomes increasingly concentrated around $\omega=\pi$.
[…]
We can also evaluate the stability of the TC phase against interactions by examining how the absolute value of the magnetization, $|m^{z}|$, varies over time. As shown in Fig. 4, $|m^{z}|$ remains constant over time for interaction strengths below a critical value $\lambda_{c}$, indicating that the TC phase is stable in this regime. The critical value $\lambda_{c}$ is dependent on the system size; for a chain of length $L=12$, it is approximately $\lambda_{c}\simeq 0.4$.
Figure 4: The absolute value of the magnetization $m^{z}$ for a chain of length $L=12$ for different values of $\lambda$, with a coupling constant $J=5.5$ and an external field $h=0.3$. The TC remains stable for interaction strengths that are weaker than $\lambda_{c}\simeq 0.4$.
III.2 Stability of the TC phase under imperfect rotations
We now examine the stability of the TC phase observed in the ITF model under conditions of imperfect rotations. Instead of executing a perfect spin-flip rotation (i.e., a rotation around the $x$-axis by an angle of $\frac{\pi}{2}$), we implement an imperfect rotation characterized by an angle $\theta=\frac{\pi}{2}-\epsilon$, where $\epsilon$ is a small positive value ($\epsilon<1$). Consequently, the time evolution operator is expressed as in Eq. (2) with the time parameters set to $t_{2}=1$ and $t_{1}=\frac{\pi}{2}-\epsilon$.
The time-dependent behavior of the magnetization along the $z$-direction is depicted in Fig. 5. In the presence of the imperfection parameter $\epsilon$, we observe that the magnetization continues to exhibit oscillatory behavior with a period of $2T$, which is twice the intrinsic period of the Hamiltonian. As $\epsilon$ increases, the amplitude of the magnetization oscillations decreases, as indicated by the declining peak value in the FFT of the magnetization $m^{z}$ (see inset of Fig. 5). Despite this reduction, the magnetization oscillations remain stable over time. However, further increases in $\epsilon$ lead to a continued decline in oscillation amplitude, ultimately destabilizing the TC phase. For a chain length of $L=12$ (refer to Fig. 5), the critical value for stable magnetization oscillations is approximately $\epsilon_{c}\approx 0.3$; beyond this point, the oscillations become unstable and their amplitude diminishes. Notably, as the chain length increases, the stability of the TC phase improves, suggesting that the critical value of $\epsilon$ may rise for larger systems. Establishing the exact value of critical $\epsilon$ will necessitate further computational simulations on larger chains, which falls outside the scope of this study.
Figure 5: The time evolution of the magnetization along the $z$-direction for various values of $\epsilon$ in a chain of length $L=12$ with coupling constant $J=5.5$ and external field $h=0.3$. The TC remains stable against a finite degree of imperfect rotation. The inset shows the FFT of the magnetization. As imperfection increases, the peak at $\omega=\pi$ diminishes and ultimately vanishes for larger $\epsilon$, signaling the breakdown of the TC phase.
IV Time Quasicrystal (TQC): general definition
In this section, we delve into the disordered ITF chain influenced by a quasiperiodic driving field, showcasing its ability to exhibit a TQC phase. Before proceeding, we will first define the TQC phase and outline the conditions necessary for the emergence of a TQC within a model’s phase diagram.
Let us consider a time-dependent Hamiltonian $\hat{H}(t)$ as followsr16 :
$\hat{H}(t)=\sum_{\vec{n}}\hat{H}_{\vec{n}}\exp(i\vec{n}\cdot\vec{\omega}t),$ (5)
where $\hat{H}_{\vec{n}}$ represents a time-independent operator, $\vec{\omega}$ is a vector of dimension $d$ with components $\omega_{1},\dots,\omega_{d}$, corresponding to the oscillation frequencies of different terms of the Hamiltonian, and $\vec{n}$ is a $d$-dimensional integer vector with components $n_{1},\dots,n_{d}$. The summation is carried out over all integer values of $n_{j}$.
When the frequencies $\omega_{j}$ in the Hamiltonian $H(t)$ are rationally related, the system exhibits periodic behavior, meaning that the dynamics of the system will repeat after a certain period $T$, which can be expressed as the least common multiple of the periods associated with the individual frequencies. In contrast, when the frequencies $\omega_{j}$ are irrationally related, the Hamiltonian becomes quasiperiodic.
Quasiperiodicity implies that the system does not repeat in a simple manner; instead, its state evolves over time in a way that covers a multi-dimensional toroidal space without returning to a previous configuration.
In the context of periodically and quasiperiodically driven systems, the concepts of time translation symmetry breaking (TTSB) and quasiperiodic time translation symmetry breaking (QTTSB) provide a framework for analyzing the behavior of operators over time.
For a periodically driven system with period $T$, TTSB occurs when the ensemble average of an operator $\hat{O}$ satisfies the condition $\langle\hat{O}(t+mT)\rangle=\langle\hat{O}(t)\rangle$ for integers $m\geq 2$. This indicates that the system exhibits a form of symmetry breaking where the observable properties of the system change over time in a different repeating manner.
On the other hand, in quasiperiodic systems, there is no single characteristic period due to the nature of their driving, which is typically defined by incommensurate frequencies. However, QTTSB can still be characterized through the frequency analysis of the system’s dynamics. Specifically, a quasiperiodic system demonstrates TQC behavior if there exists an operator $\hat{O}(t)$ such that its ensemble average can be expressed asr16 :
$\langle\hat{O}(t)\rangle=\sum_{\vec{n}}\hat{O}_{\vec{n}}\exp(i\sum_{i}\frac{n_{j}\omega_{j}}{p_{j}}t),$ (6)
where $\vec{n}$ are vectors of non-negative integers, $\omega_{j}$ are the driving frequencies, and $p_{j}$ are integers with at least one $p_{i}$ exceeding 1. This implies that the system exhibits a quasicrystalline ordering in time, characterized by multiple frequencies that do not lead to a simple periodic repetition but rather a complex structure reminiscent of quasicrystals in spatial dimensions.
In summary, TTSB refers to a breaking of periodic time symmetry in periodically driven systems, while QTTSB captures a similar notion for quasiperiodic systems through a more complex interplay of frequencies, indicating a richer dynamical structureFTS .
We now investigate the emergence of a TQC phase within the disordered ITF chain. A promising approach involves replacing the periodically driven $\pi$-pulses with a quasiperiodic driving scheme. In this context, we examine the following time-dependent Hamiltonian:
$\hat{H}(t)=\hat{H}_{ITF}(t)+\hat{H}_{drive}(t)+\frac{\Omega}{2}\sum_{i}\hat{\sigma}_{i}^{z}$ (7)
where,
$\hat{H}_{\text{ITF}}(t)=h_{1}(t)\sum\limits_{i}[-J_{i}\hat{\sigma}_{i}^{z}\hat{\sigma}_{i+1}^{z}+h_{i}\hat{\sigma}_{i}^{x}],$ (8)
and
$\hat{H}_{drive}(t)=h_{2}(t)\sum\limits_{i}[\cos(\Omega t)\hat{\sigma}_{i}^{y}-\sin(\Omega t)\hat{\sigma}_{i}^{x}].$ (9)
The time-dependent Hamiltonian in Eq. (7) consists of three main terms: (1) ITF Hamiltonian $\hat{H}_{ITF}$ which describes the interactions in the system, including nearest-neighbor couplings and a transverse field. (2) Driving Hamiltonian $\hat{H}_{drive}$ which introduces a time-dependent driving force that oscillates sinusoidally. The driving terms involve both the $x$ and $y$-components of the spin operators, which can induce interesting dynamics in the system. (3) Additional magnetic field term $\frac{\Omega}{2}\sigma_{i}^{z}$, representing a static magnetic field in the $z$-direction, which can have a significant impact on the dynamics of the spins.
The time-dependent coefficients $h_{1}(t)$ and $h_{2}(t)$ are activated and deactivated periodically allowing for a controlled temporal evolution of the system. The definitions of $h_{1}(t)$ and $h_{2}(t)$ are critical for controlling the dynamics of the system. The piece wise definition ensures that the effects of the ITF and the driving terms do not overlap in time, allowing for distinct and potentially non-trivial time evolution dynamics. This can lead to a rich phase structure, especially when considering the interplay between disorder and driving.
We define $h_{1}(t)$ and $h_{2}(t)$ as follows:
$h_{1}(t)=\begin{cases}0,&\quad\frac{T}{4}>t\quad\text{and}\quad t<\frac{-T}{4}\\ 1,&\quad\frac{-T}{4}\leq t\leq\frac{T}{4}\end{cases}$ (10)
$h_{2}(t)=\begin{cases}\frac{1}{2},&\quad\frac{T}{4}>t\quad\text{and}\quad t<\frac{-T}{4}\\ 0,&\quad\frac{-T}{4}\leq t\leq\frac{T}{4}\end{cases}$
The schematic provided in FIG. 6 likely illustrates how the system transitions between the different Hamiltonian regimes over time. This visual representation can be very helpful for understanding the sequence of events that govern the system’s dynamics during the driving protocol.
Figure 6: An illustration of time intervals for the effects of various Hamiltonian terms.
The driving terms $h_{2}(t)\cos(\Omega t)$ and $h_{2}(t)\sin(\Omega t)$ in Eq. (9), are consistent with the definition of quasiperiodicity as specified in Eq. 5. This consistency can be confirmed through the Fourier expansion given by:
$h_{2}(t)\cos(\Omega t)=\sum_{n_{1}=-\infty}^{\infty}\sum_{n_{2}=\pm 1}c_{n_{1}}e^{i(n_{1}\omega+n_{2}\Omega)t},$ (11)
where, $\omega=\frac{2\pi}{T}$ and $c_{n_{1}}$ are the coefficients of the expansion. The Fourier expansion of the driving terms indicates that they can exhibit quasiperiodic behavior, which is essential for the emergence of TQCs. The coefficients $c_{n_{1}}$ in the expansion will determine the specific nature of the quasiperiodicity, and the interplay between the various frequencies $n_{1}\omega+n_{2}\Omega$ can lead to complex dynamics.
V TQC in the disordered ITF model
The phenomenon of QTTSB can be characterized through the temporal evolution of the magnetizations $m^{x,y,z}=\langle\hat{\sigma}^{x,y,z}\rangle$. While quasienergy level statistics are typically employed to investigate the thermalization phase transition in many-body systems, as noted in reference r8 , our quasiperiodically driven system presents challenges in defining the Floquet operator and quasienergies. Consequently, we focus on analyzing the time evolution of the magnetizations.
[…]
The evolution of the magnetizations $m^{x,y,z}$ is initiated from the initial state $|\psi_{0}\rangle=\otimes_{i}^{L}[|\uparrow_{i}\rangle+|\downarrow_{i}\rangle]$. By averaging over 200 random realizations, we obtain the time dependence of the magnetizations $m^{x,y,z}$ over a time window of $100T$. The results, depicted in FIG. 7, reveal that when $\bar{J}\gg\bar{h}$, the subharmonic response is evident in the Fourier decomposition of the drive as expressed in Eq. (11). This response distinctly indicates QTTSB, characterized by sharp peaks at $k\frac{\omega}{2}\pm\Omega$ for odd integer $k$.
Figure 7: Top: Time evolution of the magnetizations $m^{x,y,z}$, for a chain of length $L=10$, with exchange coupling $J=5.5$ and transverse field $h=0.3$, revealing periodic oscillations in all three magnetization components. Bottom: FFT of the magnetization $m^{z}$ as a function of frequency $\omega$ showcasing distinct sharp peaks at $k\omega/2\pm\Omega$ (where $k$ is an odd integer), indicative of QTTSB and the formation of a QTC in the system.
It is important to note that in certain systems, such as the Ising model under a longitudinal magnetic field, a transformation exists that can map a quasiperiodic Hamiltonian to a periodic one, enabling the transition from TQC to TC through a unitary transformation FTS . However, in the ITF model, such a transformation is not readily detectable. Specifically, the transformation must convert a quasiperiodic drive into a periodic one while leaving the time-independent Ising Hamiltonian unchanged. While this is achievable for the Ising model in a longitudinal field, identifying a similar transformation for the ITF chain poses significant challenges. As we will demonstrate in Section VI, although the ITF chain with couplings $J_{i}\in[-J/2,J/2]$ exhibits a TC phase under a periodic drive, it fails to exhibit a TQC phase under the quasiperiodic drive described in Eq. (9).
V.1 Stability of the TQC phase in the ITF chain
In this section, we aim to systematically investigate the stability of the observed TQC phase within the disordered ITF chain, with a specific focus on the effects of interactions and imperfections, while also assessing the potential role of finite size effects on this TQC phase.
To rigorously confirm that the observed TQC phase in the ITF chain is not merely a finite size effect, it is crucial to assess the stability of this phase as the system size increases. Our exact diagonalization results for chains of lengths 4, 6, 8, and 10 demonstrate a progressive stabilization in the oscillatory behavior of the magnetization amplitude, indicating that larger system sizes lead to more stable oscillations. However, simulating larger chains poses significant computational challenges, particularly when a driving term, like that described by Eq. (9), complicates the dynamics. In this context, utilizing tensor network techniques, such as the time-evolving block decimation (TEBD) algorithm, proves advantageous. This method allows for efficient representation and manipulation of quantum states, making it particularly suited for exploring quantum many-body systems, especially in scenarios where traditional methods falter due to the exponential growth of the Hilbert space.
Furthermore, when the external field $h$ is significantly weaker than the coupling constant $J$, the system enters an MBL phase, characterized by the absence of thermalization and the presence of localized eigenstates. These traits contribute to the stabilization of the quasicrystalline phase as the system size increases. The robustness of the MBL phase against perturbations suggests that a similar resilience may also apply to the TQC phase observed in the ITF chain.
V.2 The stability of the TQC phase under interaction perturbations
We now investigate the stability of the TQC phase in the presence of interaction perturbations. Following a similar approach as before, we incorporate the interacting $\lambda$ term into the ITF Hamiltonian from Eq. (7) as:
$\hat{H}_{\text{ITF}}(t)=h_{1}(t)\sum\limits_{i}[-J_{i}\hat{\sigma}_{i}^{z}\hat{\sigma}_{i+1}^{z}+h_{i}\hat{\sigma}_{i}^{x}+\lambda\hat{\sigma}_{i}^{x}\hat{\sigma}_{i+1}^{x}].$ (12)
Given that the time-dependent coefficient $h_{1}(t)$ is periodically turned on and off within the time interval specified in Eq. (10), the interaction perturbation is similarly introduced periodically into the system during this interval.
The FFT analysis of the magnetization $m^{z}(t)$ is presented in Fig. 8 for different values of the interaction strength $\lambda$. The prominent peaks observed at $k\omega/2\pm\Omega$, where $k$ is an odd integer, suggest that the TQC phase demonstrates resilience to a range of interaction strengths.
Figure 8: Up: FFT analysis of the magnetization $m^{z}$ in the TQC phase for various values of the interaction strength $\lambda$ (up). The data, derived from exact diagonalization of the Hamiltonian for a disordered ITF chain of length 8 with parameters $J=5.5$ and $h=0.3$, are averaged over 200 random realizations. FFTs are computed over a time window of $100T$. Observed peaks at $k\omega/2\pm\Omega$ (where $k$ is an odd integer), demonstrate the resilience of TQC against interaction perturbations. Notably, as $\lambda$ increases, peak amplitudes decrease, ultimately vanishing at a critical value $\lambda_{c}$. Down: Variation of $\Delta A$ across different time windows ($\tau$) for various interaction perturbations. For a chain length of $L=8$, the TQC phase becomes unstable for interactions exceeding $\lambda_{c}=0.03$, as indicated by the observed decay at $\lambda>\lambda_{c}$. Error bars represent the standard deviation of the data, calculated across different time windows.
Two key observations emerge from the FFT plot as a function of frequency across varying interaction strengths. Firstly, the positions of the peaks remain unaffected by the interaction, indicating that the fundamental frequencies associated with the magnetization oscillations are intrinsic to the system. This invariance implies that the overall periodic behavior of magnetization is robust against perturbations introduced by the interactions. It indicates that the system retains its underlying oscillatory nature, reflecting the translational symmetry of the system.
Secondly, the variation in the amplitude of the magnetization oscillations, illustrated by changes in the peak magnitudes, indicates that while the system’s fundamental frequencies remain constant, the strength of the oscillations is influenced by interaction strength. This could suggest that stronger interactions suppress the magnetization dynamics, affecting how pronounced the oscillations are.
The lack of smaller peaks at higher frequencies suggests that these frequencies are no longer representative of the system’s dynamics. This can be interpreted in various ways: it may signify a damping effect due to heightened interactions, or it could indicate a transformation in the system’s energy landscape, rendering higher frequency modes less significant. However, given that the amplitude of magnetization oscillations diminishes over time in the presence of interaction perturbations, the absence of smaller peaks at higher frequencies strongly points to a damping effect.
As the parameter $\lambda$ increases, the peak amplitudes diminish and ultimately vanish at a critical value, $\lambda_{c}$. To determine $\lambda_{c}$, we employ the short-time Fourier transform (STFT) on the magnetization $m^{z}$. We segment the time interval of $100T$ into ten equal windows, performing a Fourier transform within each window to extract the amplitude of the first peak for varying values of $\lambda$. The difference in amplitude of the first peak at a non vanishing $\lambda$ and the amplitude at $\lambda=0$, denoted as $\Delta A$, provides valuable insight into the critical value $\lambda_{c}$. As shown in Fig. 8-(down), $\Delta A$ approaches zero for $\lambda$ values below $\lambda_{c}$, indicating that the TQC remains stable against interactions weaker than $\lambda_{c}$. For a chain of length $L=8$, the critical interaction value is approximately $0.03$, as illustrated in Fig. 8-(down).
The TQC phase is not merely a finite-size effect; it manifests in both finite disordered ITF chains and infinite chains alike. To assess the stability against interactions, we examined a chain of length $L=8$. It is evident that since the system resides in the MBL phase for $h<J$, we anticipate that the stability of the TQC phase will improve with increasing system size.
V.3 The stability of the TQC phase under imperfect rotations
We now examine the stability of TQC under imperfect rotations. To this end, we modify the drive term of the Hamiltonian in Eq. (7) as:
$\hat{H}_{drive}(t)=h_{2}(t)\sum\limits_{i}[(1-\epsilon)\cos(\Omega t)\hat{\sigma}_{i}^{y}-\sin(\Omega t)\hat{\sigma}_{i}^{x}],$ (13)
where $\epsilon$ cause imperfect rotation of spins. The modified drive term introduces a scaling factor $(1-\epsilon)$ to the $\cos(\Omega t)$ term, which captures the effect of imperfect rotations. The FFT of the magnetization $m^{z}$ is illustrated in Fig.9 for various values of the parameter $\epsilon$. The impact of imperfection is similar to the interaction perturbation. A key observation from the FFT of the magnetization $m^{z}$ presented in Fig. 9 is that the presence of imperfections leads to a reduction in the amplitude of the subharmonic peaks without affecting their positions. Specifically, these subharmonic peaks appear at locations given by $k\frac{\omega}{2}\pm\Omega$ where $k$ is an odd integer, indicating that while the system’s frequency structure remains stable, the strength of the response diminishes as $\epsilon$ increases.
Figure 9: Up: FFT analysis of the magnetization $m^{z}$ in the TQC phase for various values of the imperfection $\epsilon$. The data, derived from exact diagonalization of the Hamiltonian for a disordered ITF chain of length 8 with parameters $J=5.5$ and $h=0.3$, are averaged over 200 random realizations. FFTs are computed over a time window of $100T$. Observed peaks at $k\omega/2\pm\Omega$ (where $k$ is an odd integer), demonstrating the resilience of TQC against imperfections. Notably, as $\epsilon$ increases, peak amplitudes decrease, ultimately vanishing at a critical value $\epsilon_{c}$ ($\sim 0.01$ for a chain of length 8). Down: Variation of $\Delta A$ across different time windows ($\tau$) for various imperfections. For a chain length of 8, the TQC phase becomes unstable for imperfections exceeding $\epsilon_{c}=0.01$, as indicated by the observed decay at $\epsilon>\epsilon_{c}$.
This stability of the peak locations suggests that our system can tolerate a certain degree of imperfections in the rotation, represented by $\epsilon$, without losing the essential characteristics that lead to a TQC phase. As such, even in the presence of imperfect rotations, the system retains its features and can continue to function robustly as a TQC.
To determine $\epsilon_{c}$, we study the behavior of the difference in the amplitude of the first peak at a non vanishing $\epsilon$ and the amplitude at $\epsilon=0$, denoted as $\Delta A=A(\epsilon\neq 0)-A(\epsilon=0)$, for a chain of length 8. As shown in Fig. 9-(down), $\Delta A$ approaches zero for $\epsilon$s below $\epsilon_{c}=0.01$, indicating that the TQC remains stable against imperfections smaller than $0.01$.
In summary, the findings indicate that the TQC phase can be maintained under the influence of rotation imperfections, provided that these imperfections do not exceed a certain threshold. This resilience is a promising aspect of implementing TQC in practical scenarios, where imperfections are often inevitable.
As demonstrated, the critical perturbation thresholds for stability in TCs are significantly higher, by an order of magnitude, than those observed in TQCs, indicating that TCs exhibit markedly greater robustness to perturbations. This enhanced robustness is attributed to the phenomenon of quantum frequency lockingr27 , which occurs when a quantum system with an inherent characteristic frequency $\omega_{c}$ is driven at a frequency $\Omega=2\pi/T$ that is close to a rational multiple $q/r$ of $\omega_{c}$. Under these conditions, the system can achieve stable oscillations with a period precisely aligned to an integer multiple of the driving period, $qT$. In contrast, TQCs do not benefit from this frequency locking, leading to magnetization characterized by incommensurate oscillations.
VI Absence of TQC phase in ITF chain with $J_{i}$ in symmetric intervals
In this section, we investigate the absence of the TQC phase in the ITF model described by Eq. (7), where the exchange couplings $J_{i}$ are selected from a symmetric interval of $[-J/2,J/2]$. This selection implies that the exchange interactions can take both positive and negative values, leading to a rich landscape of magnetic behavior.
For weak magnetic field strengths ($h\ll J$), the randomness in the signs of $J_{i}$ plays a crucial role in determining the magnetic properties of the system. Specifically, it has been shown that both the ground state and all excited states exhibit zero net magnetizationr20 . This phenomenon arises because the competing interactions from the random couplings effectively cancel each other’s contributions to the overall magnetization.
Despite the absence of net magnetization, the system is not in a trivial paramagnetic phase; instead, it resides in a spin glass (SG) phaser24 ; r25 ; r26 . This conclusion is supported by the observation that the Edwards-Anderson order parameter is non-zero EA , indicating the presence of localized magnetic order amidst the disordered couplings. In a spin glass phase, the system can sustain a finite density of domain walls, which become immobilized due to the underlying disorder. This immobilization results in locally broken symmetry, as different regions of the system may adopt distinct magnetic configurationsr23 ; D.Huse .
As a result, the eigenstates of the Hamiltonian are characterized by random glassy configurations, showcasing a complex interplay between disorder and magnetic interactions. These configurations reflect the frustration and variability inherent in the system, leading to a diverse array of possible states that can be accessed under different conditions. This behavior is characteristic of spin glasses, where the combination of randomness and interaction leads to a non-trivial energy landscape, challenging conventional notions of order in magnetic systems.
Consider a chain of length $L$ featuring the emergence of two domain walls located at sites $j_{1}$ and $j_{2}$. The quantum state corresponding to this configuration can be expressed in one of the following forms:
$|j_{1},j_{2}\rangle=|\sigma_{1},\dots,\sigma_{j_{1}-1},j_{1},\sigma_{j_{1}+1},\dots,\sigma_{j_{2}-1},j_{2},\sigma_{j_{2}+1},\dots,\sigma_{L}\rangle,$
$|\bar{j}_{1},\bar{j}_{2}\rangle=|\bar{\sigma}_{1},\dots,\bar{\sigma}_{j_{1}-1},j_{1},\bar{\sigma}_{j_{1}+1},\dots\bar{\sigma}_{j_{2}-1},j_{2},\bar{\sigma}_{j_{2}+1},\dots,\bar{\sigma}_{L}\rangle,$
where, each spin variable $\sigma_{i}$ can randomly take values of $+1$ or $-1$, with the constraint that $\bar{\sigma}=-\sigma$. Under the influence of magnetic fields, the system experiences tunneling between the states $|j_{1},j_{2}\rangle$ and $|\bar{j}_{1},\bar{j}_{2}\rangle$. As a result, the Hamiltonian’s eigenstate emerges as a coherent superposition of these states as:
|$|\{j_{1},j_{2}\},p=\pm 1\rangle=\frac{1}{\sqrt{2}}(|j_{1},j_{2}\rangle\pm|\bar{j}_{1},\bar{j}_{2}\rangle).$ (14)
These states, characterized by the expectation values of the parity operator $\hat{P}$ and the Ising-symmetric domain wall operators $\hat{D}_{j_{1}}\equiv\hat{\sigma}^{z}_{j_{1}}\hat{\sigma}^{z}_{j_{1}+1}$ and $\hat{D}_{j_{2}}\equiv\hat{\sigma}^{z}_{j_{2}}\hat{\sigma}^{z}_{j_{2}+1}$, exhibit Ising symmetry breaking and long-range order, as evidenced by the non-zero spin-spin correlators in each eigenstate:
$C_{ij}=\langle\{j_{1},j_{2}\},p|\sigma_{i}^{z}\sigma_{j}^{z}|\{j_{1},j_{2}\},p\rangle,~{}~{}\lim_{|i-j|\rightarrow a.l.}C_{ij}\neq 0,$ (15)
where ’$a.l.$’ denotes arbitrarily large distances.
The driven ITF model exhibits rich phases, particularly when compared to its undriven counterpart. In the undriven disordered ITF model, the primary phases observed are the PM and SG phases. However, the introduction of driving field leads to the emergence of additional dynamical phases.
In the ITF model periodically driven by an external magnetic field as in Eq. (1), the two notable dynamical phases arer19 ; r7 :
1. $\pi$PM-MBL phase: This phase is characterized by a paramagnetic behavior under the influence of a periodic driving field, where the system exhibits localization properties typical of MBL phases. In this regime, the quantum coherence and entanglement are preserved, despite the presence of disorder and external driving.
2. $\pi$SG-MBL phase: In this phase, the system displays spin glass characteristics while still retaining MBL traits. The dynamics are influenced by the interplay between the disorder, the transverse field, and the driving, leading to a scenario where the system can remain localized while exhibiting complex spin configurations typical of spin glass states.
In the periodically driven spin glass phase, our findings (not shown) demonstrate the stability of the TC phase even amidst interaction perturbations and imperfect rotations. The system persists in the localized phase for interaction strengths ranging from $0<\lambda\leq 1$ r7 , highlighting its resilience to these perturbations while maintaining its position in the TC phase.
[…]
Our investigations into the time evolution of the magnetizations $m^{x,y,z}$ indicate a notable absence of TQC in either the SG-MBL or the PM-MBL phases under the current configuration. This observation is in stark contrast to scenarios where $J_{i}$ is selected from the range $[J/2,3J/2]$, where quasiperiodic dynamics can manifest. The lack of a TQC in the present setup can be primarily attributed to the characteristics inherent to the spin glass phase. The presence of domain walls disrupts the potential for the system to exhibit dynamics that repeat in a quasiperiodic fashion over time. Consequently, there are no sustained oscillations in magnetization with a robust frequency, underscoring the intricate interplay between disorder and driving in shaping the dynamical properties of spin systems.
This highlights the complexity of quantum systems under external influences and the significant role that disorder plays in determining their dynamical behavior, ultimately uncovering a complex interplay of phase interactions that deserves deeper investigation.
VII Conclusion
In summary, our investigation provides a thorough understanding of the dynamical properties of the disordered Ising model in a transverse magnetic field (ITF) under periodic and quasiperiodic driving conditions.
In the first part of our study, we demonstrated that the driven disordered ITF chain exhibits time-dependent magnetization $m^{z}$ characterized by significant oscillations with a period of $2T$ and an oscillatory-decaying envelope. This stability against various perturbations highlights the robustness of the time crystal (TC) phase, indicating its potential applications in quantum technologies.
In the second part, we explored the ability of the disordered ITF chain to realize a time quasicrystalline (TQC) phase. By applying a sinusoidal driving field with frequencies that are irrationally related, we observed complex frequency structures in the magnetization that resemble those of quasicrystals. This finding paves the way for further exploration of TQC in disordered systems, showcasing the rich dynamical behavior that can arise from quasiperiodic driving.
Our work establishes a foundation for future experimental implementations in platforms such as cold atomic systems and quantum simulators. The insights obtained from TQCs hold significant implications for quantum information, quantum computing, and the development of novel materials with specific dynamical responses. The interplay between disorder, driving, and quantum coherence in these systems calls for further investigation, potentially leading to breakthroughs in understanding and harnessing complex quantum behaviors.
VIII Acknowledgment
We would like to express our gratitude to Seyed Saeed Soyouf Jahormi for his insightful discussions. Additionally, JA extends his appreciation to the Institute for Advanced Studies in Basic Sciences for their financial support through research Grant No. GIASBS12969.